# Do Multimodal Large Language Models See Like Humans?

### v1

@@ 4.3 Ablation Study
Model Size. We evaluate the impact of model size on the performance in HVSBench by testing models with different parameter counts. For this study, we select two representative methods: DeepSeek-VL [28] and GPT-4o [1], from both open-source and proprietary MLLMs, to provide a comprehensive analysis. As shown in Table 4, larger MLLMs generally outperform smaller ones across all metrics. It suggests that increasing model size leads to better alignment with HVS for MLLMs.
@@
Human Captions or Descriptions Improve Alignment? It is possible that certain annotations, like human-generated captions and detailed descriptions, implicitly reflect the HVS. Therefore, we evaluated whether adding detailed descriptions (“Detail.”) or short captions (“Cap.”) could improve the performance of the base GPT-4o model.
@@
Field-Specific Hints Improve Alignment? Since definitions and prior knowledge of the fields also provide cues for the HVS, we examine whether adding Field-Specific Hints can enhance model performance. For example, the hint for the Prominence field is: The detection of salient objects aims to simulate the human visual perception system by identifying and localizing the most visually striking object(s) in a scene. Previous research suggests that the most salient object is the one that attracts the highest proportion of fixations, as indicated by the agreement between fixation patterns and saliency judgments. In general, there are two primary priors: objects closer to the viewer are perceived as more salient, and salient objects often appear near the center of the scene. Additionally, cues such as color contrast, spatial bias, and depth contrast also influence saliency. Refer to Sec. E in supplementary for Field-Specific Hints in each field.
@@
4.4 Possible Applications
We design a Cropping-Based Prominence Enhancement, aiming to show one of the applications by a better HVS-aligned MLLM in the prominent field. Specifically, we examine how GPT-4o crops the image to enhance the prominence of one object: a photo. GPT-4o with a task-specific hint generates a reasonable analysis and successfully crops the image to highlight the photo, compared to the result without hint, demonstrating better alignment with HVS. This can be directly applied to automated design, context-aware content generation, and visual storytelling.
Such improvements in alignment with HVS are crucial, not only for enhancing the performance of vision models [41], but also for facilitating emerging applications with broad societal and technological impact, e.g., perception in autonomous vehicles and robots, human-robot collaboration, assistive tools for the visually impaired, and even tasks like saliency ranking [10].
Figure 7: One of our applications: Prominence Enhancement.
@@ 5 Conclusion
    Section F: Detailed Benchmark Results
    This section contains a detailed report of benchmark results for all 13 question types, offering a comprehensive view of the performance across various fields and question types.
@@ Appendix C Detailed Settings
Note that we allocate 10% of HVSBench for evaluation purposes, while reserving the remaining 90% for further exploration or applications, such as instruction-tuning of MLLMs.
@@ 2.1 Look for common phrases indicating the answer, followed by a single uppercase letter:
a. Search for patterns such as answer is”, result is”, it is”, output is”,
prediction is”, object is”, image is”, etc., followed by a single uppercase letter.
@@
iii. If no letter is found, append an empty string (“”) to cleaned_labels.
@@ 2.1 Search for specific patterns indicating a list of answers:
a. Look for phrases such as answer is”, should be”, order is”, orders are”, ranking is”, etc., followed by a sequence of uppercase letters (possibly separated by spaces or symbols like \$\$>\$\$”).
@@ 2.2 If no specific pattern is matched:
2.3 If no uppercase letters are found in the item, append an empty string (“”).
@@
Pseudo-code 11 shows our standardization for Q3 in the Prioritizing field.

### v2

@@ 4.3 Ablation Study
Model Size. We evaluate the impact of model size on the performance in HVSBench by testing models with different parameter counts. For this study, we select two representative methods: DeepSeek-VL [34] and GPT-4o [1], from both open-source and proprietary MLLMs, to provide a comprehensive analysis. As shown in Table 4, larger MLLMs generally outperform smaller ones across all metrics. It suggests that increasing model size leads to better alignment with HVS for MLLMs. Refer to the supplementary for more experiments on this aspect.
@@
Human Captions or Descriptions Improve Alignment? It is possible that certain annotations, like human captions from COCO caption and detailed descriptions from LLaVA-Instruct-150K, implicitly reflect the HVS. Therefore, we evaluated whether adding detailed descriptions (“Detail.”) or short captions (“Cap.”) could improve the performance of the base GPT-4o model.
@@
Field-Specific Hints Improve Alignment? Since definitions and prior knowledge of the fields also provide cues for the HVS, we examine whether adding Field-Specific Hints can enhance model performance. For example, the simplified hint for the Prominence field is: The detection of salient objects aims to simulate the human visual perception system by identifying and localizing the most visually striking object(s) in a scene. … (omitted), cues such as color contrast, spatial bias, and depth contrast also influence saliency. Refer to Sec. E in supplementary for full hints in each field.
@@
4.4 Discussion: Benefits and Applications
While human perception mimicry may not be universally beneficial, but we stress its value in key tasks. (1) Mimicking human perception can significantly improve tasks requiring precise visual grounding, such as QA and captioning. In table 4 of [53], models mimicking human fixation mechanisms achieve shorter search lengths in QA tasks, improving efficiency and accuracy. (2) Content generation models better aligned with the HVS can produce more reasonable outputs. Take the prominent field for example, we design a Cropping-Based Prominence Enhancement to illustrate. Specifically, we examine how GPT-4o crops the image to enhance the prominence of one object: a photo. GPT-4o with a task-specific hint generates a reasonable analysis and successfully crops the image to highlight the photo, compared to the result without hint, demonstrating better alignment with HVS. This can be directly applied to automated design, context-aware content generation, and visual storytelling.
Figure 7: One generation application: Prominence Enhancement.
(3) In HVS-related downstream tasks, models better aligned with HVS perform better. For example, Qwen2-VL outperforms MiniCPM-V in our HVSBench and also in general tasks such as human perception, visual illusion (MMMU, HallusionBench) , but is comparable or worse in tasks unrelated to HVS like math reasoning (MathVision). (4) Practical applications like autonomous driving, perception in robots, assistive tools for the visually impaired, and tasks directly related to the HVS (e.g., saliency ranking), further demonstrate the benefits where HVS alignment ensures intuitive, user-centered outputs. Based on these four aspects, we believe mimicking human perception will benefit machine vision.
@@ 5 Conclusion
    Section F: Detailed Benchmark Results and Disucussions
    This section contains a detailed report of benchmark results for all 13 question types, offering a comprehensive view of the performance across various fields and question types. We also provide additional ablation studies and related discussions for them.
  7. 7.
    Section G: More Related Work
    We include more related work for different research areas to showcase the necessity of our HVSBench.
@@ 2.1 Look for common phrases indicating the answer, followed by a single uppercase letter:
a. Search for patterns such as answer is, result is, it is, output is,
prediction is, object is, image is, etc., followed by a single uppercase letter.
@@
iii. If no letter is found, append an empty string (“) to cleaned_labels.
@@ 2.1 Search for specific patterns indicating a list of answers:
a. Look for phrases such as answer is, should be, order is, orders are, ranking is, etc., followed by a sequence of uppercase letters (possibly separated by spaces or symbols like
@@ 2.2 If no specific pattern is matched:
2.3 If no uppercase letters are found in the item, append an empty string (“).
@@
Pseudo-code C shows our standardization for Q3 in the Prioritizing field.
@@ Appendix F Detailed Benchmark Results
F.1 Ablation on Model Size
We evaluate the impact of model size on the performance in HVSBench by testing models with different parameter counts. For this study, we select two representative methods: DeepSeek-VL [34] and GPT-4o [1], from both open-source and proprietary MLLMs, to provide a comprehensive analysis. As shown in Table 4, larger MLLMs generally outperform smaller ones across all metrics. It suggests that increasing model size leads to better alignment with HVS for MLLMs. Refer to supplementary for more experiments.
[…]
  GPT4-o mini [1] N/A 0.3126 0.4480 0.3312 0.3560 0.3766
GPT4-o [1] N/A 0.3139 0.4512 0.3621 0.3737 0.4015
DeepSeek-VL [34] 1.3B 0.1758 0.2513 0.2950 0.3188 0.2843
7B 0.3223 0.4544 0.3327 0.3445 0.3516
Qwen2-VL [52] 2B 0.0220 0.3499 0.1100 0.2439 0.2643
7B 0.4103 0.5090 0.3901 0.3182 0.4214
72B 0.4957 0.4889 0.4158 0.4299 0.5810
F.2 Discussion: Why do MLLMs work on HVS-related tasks?
The design of MLLMs (particularly transformer-based architectures) allows approximations of human visual capabilities. For Prominence, Subitizing and Prioritizing, [44] leverages MLLMs to derive visual saliency hierarchies as the guidance for saliency prediction, proving MLLMs’ potential capacity to mimic human visual prioritization. For Free-Viewing and Searching, [56] can predict scanpaths by simulating gaze patterns using Transformer-based attention mechanisms. For Prominence, Subitizing and Prioritizing, [47] shows that human-like saliency can be simulated by attention layers inherently learning to weight salient regions of an input image, mirroring human prioritization of significant elements. We highlight the value of benchmarking HVS alignment, as it benefits many tasks.
Appendix G More Related Work